72. Currently, you owe the bank $19,600 for a car loan. The loan has an interest rate of 7.75 percent and monthly payments of $620. Your financial situation recently changed such that you can no longer afford these payments. After talking with your banker and explaining the situation, he has agreed to lower the monthly payments to $450 while keeping the interest rate at 7.75 percent. How much longer will it take you to repay this loan than you had originally planned?

73. You have $25,000 saved in a bank account paying monthly interest with an APR of 2.0%. If you deposit an additional $1,000 each month beginning in one month and ending in one year, how much will be in the account immediately after the last deposit (and assuming you didn’t take anything out!)?

74. A financial institution is selling a unique structured investment (well, not really, it’s just a corporate bond). The investment will require investors to pay $1.1M today. They will then receive ten annual cash flows of $70,000 beginning in one year. And a single cash flow of an additional $1M coinciding with the last of the $70,000 making the last payment at year ten equal $1,070,000. What annual return are investors earning on this investment?
Answer #1

72. PV of car loan =19600
Rate per month =7.75%/12
PMT initially =620
Number of Periods can be calculated using Financial calculator
I/Y =7.75%/12;PMT=620;PV=-19600;CPT N =35.4739 months

If PMT paid =450
Number of Periods can be calculated using Financial calculator
I/Y =7.75%/12;PMT=450;PV=-19600;CPT N =51.3091 months

Months longer =51.3091-35.4739 =15.84 months

73. PV =25000
Rate per month =2%/12
PMT per month =1000
Number of Periods =12
FV of amount in 1 year =PV*(1+r)^n+PMT*((1+r)^n-1)/r)=25000*(1+2%/12)^12+1000*((1+2%/12)^12-1)/(2%/12))
37615.22

74. Initial Investment or PV =-1100000
PMT for 10 years =70000
Single last paymeny at end of 10 years =1000000
Using Financial calculator to calculate annual rate of return
N=10;PMT=70000;PV=-1100000;FV=1000000;CPT I/Y =5.66%
Annual Rate of return =5.66%